OpenStudy (anonymous):

5/8 of 56 = ?? 4/9 of 81 = ?? help

OpenStudy (anonymous):

40 36 Let me explain.

OpenStudy (anonymous):

pls do explain :)

OpenStudy (anonymous):

The top one is 35. That was a blunder. Both numbers can be evenly separated by the denominators.

OpenStudy (igreen):

'Of' means Multiplication. \(\Large\frac{5}{8}\) \(\times~ 56\) \(\Large\frac{4}{9}\) \(\times~ 81\)

OpenStudy (anonymous):

The top each group will have 7. The bottom each group will have 9.

OpenStudy (igreen):

Then you can just type them in Google..and it gives you 35 and 36..

OpenStudy (anonymous):

ohh i get it :) THANKS GUYS!!

OpenStudy (anonymous):

Or multiplication as stayed by iGreen.